An entry pass to a museum costs $12 for a child for a child and $15 for an adult. A total of 300 entry passes were sold, and the total money collected is $4140. The number of $15 entry passes sold are
Answer by ankor@dixie-net.com(22740)

An entry pass to a museum costs $12 for a child for a child and $15 for an adult. A total of 300 entry passes were sold, and the total money collected is $4140. The number of $15 entry passes sold are
:
let a = no. of adult tickets
let c = no. of children's
The total tickets sold equation
a + c = 300
c = -a + 300, use this form for substitution
The $$ equation
15a + 12c = 4140
simplify, divide by 3
5a + 4c = 1380
replace c with (-a+300)
5a + 4(-a+300) = 1380
5a - 4a + 1200 = 1380
a = 1380 - 1200
a = 180 adults,($15 adult passes)
:
:
Confirm this, find the no. of children
300 - 180 = 120 children
calculate how much of each
15(180) = $2700
12(120) = $1440
----------------
total $: $4140